OpenStudy (anonymous):

Solve 5^3x + 1 = 25x + 1 a.x = 4 b.x = -2 c.x = 1 d.x = 0

OpenStudy (anonymous):

first move all your "X"'s to one side and all your integers to another: \[5^3x-25x=-1+1\] Now simply solve for "X" 5^3x-25x=0 now no matter what you do to the other side, you will still have "0" There-fore, your answer is 0

OpenStudy (campbell_st):

hard to work out the equation \[5^{3x} + 1= 25^x + 1\] but if it is.. here's a solution write the rhs in terms of 5 to a power \[5^{3x} + 1 = (5^2)^x + 1\] this gives \[5^{3x} + 1 = 5^{2x} +1\] subtract 1 from both sides \[5^{3x} = 5^{2x} \] since the bases are the same equate the powers 3x = 2x solve for x

OpenStudy (anonymous):

^ DOH! forgot all about that lol

OpenStudy (campbell_st):

hope it helped